OpenStudy (anonymous):

Alex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 5 cm and a height of 11 cm. The dimensions of the actual water tanks that will be installed in the park are 9 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm^2. Use 22/7 for pi.

OpenStudy (anonymous):

i got 40731 as my answer. i just want to know if i am right or if i did something wrong?

OpenStudy (kropot72):

Actual radius = 5 * 9 = 45 cm Actual height = 11 * 9 = 99 cm Surface area of side is given by 2 * pi * r Area of end is given by pi * r^2

OpenStudy (anonymous):

i thought the surface area of a cylinder is 2*pi*r^2 + 2*pi*r*h

OpenStudy (kropot72):

Sorry. Surface area of side is given by 2 * pi * r * h If the cylinder is closed at both ends then the area of one end is doubled.

OpenStudy (anonymous):

yeah i assume it is closed at both ends since it doesnt say otherwise. but is 40731 the correct answer? because thats what i got using the formula i just want to make sure

OpenStudy (kropot72):

You are correct (rounded to the nearest cm^2) assuming the cylinder is closed at both ends.

OpenStudy (anonymous):

thank you :) can u please give me a medal?

OpenStudy (kropot72):

You're welcome :)